# Brain states analysis of EEG predicts multiple sclerosis and mirrors disease duration and burden

### v6

@@ Abstract
Objective: The proposed method deciphers electrical brain states which as multi-dimensional cognetoms quantitatively discriminate normal from pathological patterns in an EEG.
@@ 1 Introduction
MS is a chronic, often relapsing-remitting, inflammatory disease with multiple foci of demyelination disseminated over space and time throughout the CNS. It can lead to impaired cognitive reserves and capacity (Maggi et al., 2025; Pauletti et al., 2025) disability (Cobo-Calvo et al., 2025) and early death. Northern countries show a high frequency for MS, with the highest prevalence in Canada of 0.3% (Amankwah et al., 2017).
@@
For this work we propose an extension for the concept of brain states analysis (Janoos et al., 2011; Mórocz et al., 2012; Janoos et al., 2013). Instead of the statistical principles based on hidden Markov model (HMM) and functional MRI data we developed a novel method for brain states detection and analysis of EEG signals taken from a prior cross-sectional EEG study about patients with multiple sclerosis and healthy subjects (Jouzizadeh et al., 2021).
@@ 2.1 Subjects and EEG recording
The EEG data of 88 subjects were taken from an earlier project. Participant criteria and recording sessions have been described in detail in the original publication (Jouzizadeh et al., 2021). Participants signed consent forms for all procedures and data analyses in accordance with the Declaration of Helsinki. The anonymized data of 36 subjects diagnosed with MS (23 female, 13 male, mean age (29.6 years, age range 20-53 years) and 52 healthy volunteers (25 female, 27 male, mean age 20.7 years, age range 18-31 years) were taken into consideration for this data analysis project (see tables in Appendices A and B for subject information). The mean of the Expanded Disability Status Scale (EDSS) scores for MS patients at EEG acquisition time was 1.96 (EDSS score range 0-7). The diagnosis of MS was based on the revised McDonald criteria for relapsing-remitting MS (RRMS) as used at Boualicina Hospital, Sari, Iran. The mean disease duration after confirmed diagnosis was at EEG recording time 5.1 years (range 0-18 years). Resting-state EEG was recorded for 5 minutes for each condition, eyes-open (EO) and eyes-closed (EC), using a Brainmaster Discovery24 data acquisition system and 19 scalp electrode positions at 256 Hz sampling rate.
@@ 2.2 Data preparation
Raw binary sensor data were extracted as time stream vectors from original EEG data files without further metadata or personal information. Electrode labels and anatomical positional information were discarded and played no further role during data analysis. Attention was paid to maximally preserve the intrinsic physiological information present in the data for the brain states generation phase further downstream in the processing pipeline. EEG data were taken as is without applying common manual or automated data filtering, cleansing and artifact removal techniques for noise reduction, faulty electrodes, eye blink, cardiac or muscle activity, and hence without introducing data pre-processing related artifacts.
The need for artifact removal is an ongoing debate in the field. For example, in stroke patients who undergo real-time EEG-based brain computer interface (BCI) therapies, the paradigm-related muscular activity in naturalistic movements can carry a significant informative and therapeutic value even as artifactual non-brain-derived components. In contrast, the very same components may detrimentally affect analysis of brain physiological data (McDermott et al., 2021). Another example where noise filtering can remove a mixture of both artifactual and valid signal is found in intracranial EEG studies because artifacts linked just loosely to behavior remain poorly characterized (Meisler et al., 2019). Instead the authors suggest to increase the number of observations and so to enhance statistical power.
Resting state paradigms come with a random pattern of incidental and intentional muscular activities that surface as subject specific chance noise in space and time. The data processing pipeline during the brain state detection and the later analysis phases intrinsically suppress noise components (Janoos et al., 2011; Mórocz et al., 2012) and fundamentally render the approach apt to handle clinical-grade real-life data (Janoos et al., 2013). This ‘design-specific noise robustness’ lies in the fact that brain states are generated for the entire data set and hence hold inter-individual, universally present features but suppress intra-individual periodic or permanent signal properties. This way EEG artifacts due to motor, sensory and other effects do nor linger in brain states nor lead to subgroups in the analysis. Likewise is the subsequently computed disease predictor space not sensitive to patient-specific muscle artifacts in the electrical signal.
@@ 2.3 Brain states detection (BSD)
We were inspired by the similarity of a brain states analysis method we had developed for processing rapid functional MRI data (Janoos et al., 2011; Mórocz et al., 2012) and the concept of microstates analysis for EEG data introduced by Lehmann et al (Lehmann et al., 1987). Both approaches aim to transform the complex original data into a sequence of discrete, quasi-stable meaningful patterns called brain states or microstates. It was shown that the distribution of these states play a role as biomarkers in various psychiatric and neurological conditions including MS (Koenig et al., 1999; Gschwind et al., 2016).
We define brain states, or cognetoms, as snapshots of repeatable patterns of highly correlated signals also understood as moments of distributed brain activity of a given time point (Janoos et al., 2011; Mórocz et al., 2012). In other words, states are defined as patterns of statistical dependencies or spatio-temporal correlations between brain regions measured across EEG electrodes. In this paper, a brain state is a typical configuration of spatial activity without actual knowledge of spatial coordinates of electrodes. The states serve as lowest-level elements, or level-0 cognetoms, quasi as alphabet letters or building blocks to use in future aggregation cascade models of ever increasing cognetom complexity and abstraction level.
@@
In a second move, for the temporal coding of brain states, the best match distribution of brain states is detected for each time point along the EEG data time axis with weights shown for each brain state. Each time step in the original sequence is assigned to its nearest brain state cluster, producing a simplified temporal sequence of discrete brain state codes, each associated with signal intensities across frequency bands. In other words, our approach translates the multi-channel raw EEG signal of each subject into a sequence of brain states as qualitative feature with the associated weights as numerical quantitative feature.
@@ 3.2 Specific feature combinations for MS detection
The relative importance of 80 feature combinations (8 brain states x 5 frequency bands x 2 experimental conditions) that separate the two cohorts into healthy vs MS is outlined in figure 2. We observe that a set of feature combinations of high importance include both eyes-open and eyes-closed conditions and the delta frequency band that appear to drive the separation between healthy and MS subjects. Averaged frequency spectra for both cohorts are shown as graphs in Appendix C.
@@ 3.3 Disease predictor compared to clinical data
The brain states analysis technique separates with surprising ease and high probability healthy controls from patients with MS (figures 1 and 3). Considering the data to originate from single clinical-grade EEG recordings, we therefore examined additional objective neurological measures such as EDSS values or plain disease duration which we hypothesized will correlate with our EEG analysis. These last two clinical measures arguably reflect disease related accumulation of neurological deficits.
@@
Both plots in figure 3 imply an important correlation of these two clinical measures with the brain states predictor space, which, we surmise, distinguishes the better the two cohorts, the more disease burden accumulates in the CNS. In fact, all patients to the far left in the ‘control’ brain states predictor space (marked green with letters from a to g) feature a short disease duration of less than two years (mean 0.8 years) and cited motor and sensory deficits which may originate from lesions in the spinal cord or lower brain rather than upper brain parts. In contrast, patients to the right marked red, with a mean disease duration of 6.0 years, reported also vision, balance or brain stem disturbances beside motor and sensory issues in the limbs. Still, those patients in red to the right with a brief disease duration of 3 years or less, may in fact, we hypothesize, suffer from a longer, hidden disease duration, albeit clinically compensated, to amass enough supratentorial pathology that finally deranges their EEG patterns. We assume that an increase in lesion load in the cerebrum, the principal generator for the EEG signal, relates just vaguely and non-linearly to the emerging complex clinical neurological deficits, be that in cognition, EDSS, or then in form of clinical disease duration, or ‘true’ disease onset, an unknown entity for that matter.
@@
Figure 3: Distinction of control subjects from patients with MS through brain states prediction. The disease predictor value corresponds to the rotated coordinates as seen in Fig. 1B. A rotation was performed using the principal components of the point distribution where the predictor value is considered as the primary principal component. (A) The upper graph illustrates this separability plotted against EDSS measures that were obtained in MS. Patients are shown as red or green dots while controls feature blue dots. Seven patients are found as outliers on the left side of the disease predictor axis and are marked with green characters (a-g) for further discussion under Results. A linear regression is shown as gray line. (B) The lower graph plots separability against disease duration in years. Two patients (d,f) lack data on disease duration and hence are not included here. Linear regression is shown as gray trend line. The source for the post-analysis correlation calculation is in the values for the brain state predictor space and the other two modalities such as the EDSS value or the disease duration in years. For the calculation itself the open-source python module linregress from scipy was appplied (Virtanen et al., 2020).
@@ 4 Discussion
We successfully transformed a brain states hypothesis (Janoos et al., 2011) into an analytic method for processing EEG data using a quantitative embedding and modeling approach for brain states and applied the new technique to separate the two subject cohorts of an earlier EEG study about MS (Jouzizadeh et al., 2021). We observed several beneficial aspects when brain states analysis is added in a context as described here.
@@ 4.1 Implications
First we demonstrate with an initial implementation of our novel algorithm, a two-phase processing pipeline with brain states detection and brain states analysis, that already regular clinical EEG data clearly carry the necessary information to discriminate two cohorts as shown here. This is to our knowledge a first successful demonstration of such a hypothesis-free space-agnostic data-driven full-brain raw-data EEG analysis approach for MS. Second we state that the inclusion of brain states in the analysis is sensible and enhances the statistical robustness of the approach. Third we believe this brain states approach is easily generalizable to other disease conditions and importantly also for the intra-individual monitoring of serially acquired longitudinal EEG data.
The current method readily distinguished healthy from MS cases using clinical EEG data in spite of obstacles faced with : i) early stage of the software implementation; ii) swap in the original purpose of the approach from monitoring serial data to comparing subjects; iii) low information density in common clinical EEG recordings. We explain the analysis outcome with flow derangement between electrical mental processes in MS secondary to altered physical brain properties (Kisler et al., 2020; Paolicelli et al., 2021; Salim et al., 2021; Wątorek et al., 2024). This was newly confirmed in recent studies about disturbances of conductance velocity (Sorrentino et al., 2022; Wagner et al., 2025) although we currently offer no method to quantify processing speed in brain states. However, we already know that certain feature combinations of high relevance carry slow EEG components associated with both eye-open and eyes-closed conditions, that appear to prominently drive, but still as part of the entire train of feature combinations, the separation of controls from patients (see figure 2).
@@ 4.2 Significance
That this method succeeds on clinical-grade EEG data rests on two facts, in the processing of complete brain states and the characterizing of entire EEG recordings in form of comprehensive brain states sequences without cleansing noisy epochs. These two points highlight potential advantages of the method in that older less-advanced EEG databases could be sensibly re-analyzed with modern brain states analysis techniques, while collecting new EEG data in MS or other conditions could for budgetary reasons happen on affordable low-resolution EEG recording devices. Having said that, only high sensor density (Formica et al., 2025; Wątorek et al., 2024), high scan frequency and rich cognitive stimulative task conditions will leverage dense enough information in the EEG and hence contribute to deep brain states abstraction models. While clinical EEG data come in their own right only the highest-quality EEG data can deliver the fodder to study the internal syntax inside multi-hierarchical cognetom cascade models.
@@ 4.3 Limitations
Third, the brain states epoch window kept for data-technical reasons at two seconds width proved to work surprisingly well but does no justice to the speed at which ephemeral electrical brain activities operate. In other words, it is the acquisition of high-quality EEG data that will crucially counter such limitations and properly account for the short-lived transient nature of informative high-frequency signal elements.
Fourth, we are unaware of the medication given to study participants nor was this information conveyed prior (Jouzizadeh et al., 2021). Certain drug classes are known to modify physiological brain responses where patients with MS may indeed receive periodically benzodiazepines (Van Schependom et al., 2019) for example to control muscle spasms and sleep disturbances, though such symptoms tend to occur at later disease stages. This makes a benzodiazepine induced bias in the analysis outcome of the present study less likely in view of the short mean disease duration of 5.1 years. Similarly does the predominant delta character of the leading brain states feature combinations (see figure 2) point away from the beta frequency presence in benzodiazepine treated patients with MS (Van Schependom et al., 2019).
@@ 4.4 Future objectives
First we see the purpose of this publication in the introduction of a novel hypothesis-free data-driven analysis method tested on EEG and clinical data of patients with MS. However another reason is our desire to raise interest in the community towards functional brain data in longitudinal clinical studies in MS in the form of serial EEG measures (Kiiski et al., 2018).
We envision further improvements in the brain states detection (BSD) process through widening the spectrum of methods involved in analyzing the EEG signal. Analysis through multifractality and persistence effectively characterizes EEG recordings of patients with MS during eyes open and eyes closed conditions (Wątorek et al., 2024). We consider multifractality and information about long-range temporal correlations, beside FFT spectra and microstates, as procedural candidates that enrich the brain states feature space and enhance the design for time series analyses in future phases of this project.

### v7

@@ Abstract
Objective: The proposed method deciphers electrical brain states which as multi-dimensional space-neutral cognetoms quantitatively discriminate normal from pathological patterns in an EEG.
@@ 1 Introduction
MS is a chronic, often relapsing‑remitting, inflammatory disease with multiple foci of demyelination disseminated over space and time throughout the CNS. It can lead to impaired cognitive reserves and capacity (Maggi et al., 2025; Pauletti et al., 2025) disability (Cobo-Calvo et al., 2025) and early death. Northern countries show a high frequency for MS, with the highest prevalence in Canada of 0.3% (Amankwah et al., 2017).
@@
For this work we propose an extension for the concept of brain states analysis (Janoos et al., 2011; Mórocz et al., 2012; Janoos et al., 2013). Instead of the statistical principles based on hidden Markov model (HMM) and functional MRI data we developed a novel space-neutral method for brain states detection and analysis of EEG signals taken from a prior cross-sectional EEG study about patients with multiple sclerosis and healthy subjects (Jouzizadeh et al., 2021).
@@ 2.1 Subjects and EEG recording
The EEG data of 88 subjects were taken from an earlier project. Participant criteria and recording sessions have been described in detail in the original publication (Jouzizadeh et al., 2021). Participants signed consent forms for all procedures and data analyses in accordance with the Declaration of Helsinki. The anonymized data of 36 subjects diagnosed with MS (23 female, 13 male, mean age (29.6 years, age range 20-53 years) and 52 healthy volunteers (25 female, 27 male, mean age 20.7 years, age range 18-31 years) were taken into consideration for this data analysis project (see tables in A and B for subject information). The mean of the Expanded Disability Status Scale (EDSS) scores for MS patients at EEG acquisition time was 1.96 (EDSS score range 0-7). The diagnosis of MS was based on the revised McDonald criteria for relapsing-remitting MS (RRMS) as used at Boualicina Hospital, Sari, Iran. The mean disease duration after confirmed diagnosis was at EEG recording time 5.1 years (range 0-18 years). Resting-state EEG was recorded for 5 minutes for each condition, eyes-open (EO) and eyes-closed (EC), using a Brainmaster Discovery24 data acquisition system and 19 scalp electrode positions at 256 Hz sampling rate.
@@ 2.2 Data preparation
Raw binary sensor data were extracted as time stream vectors from original EEG data files without further metadata or personal information. Electrode labels and anatomical positional information played no role during data analysis. Attention was paid to maximally preserve the intrinsic physiological information present in the data (Delorme, 2023) for the brain states generation phase further downstream in the processing pipeline. EEG data were taken as is without applying common manual or automated data filtering, cleansing and artifact removal techniques for noise reduction, faulty electrodes, eye blink, cardiac or muscle activity, and hence without introducing data pre-processing related artifacts.
@@ 2.3 Brain states detection (BSD)
We were inspired by the similarity of the brain states analysys method we had developed for functional MRI data (Janoos et al., 2011; Mórocz et al., 2012) and the concept of microstates analysis for EEG data introduced by Lehmann et al (Lehmann et al., 1987). Both approaches aim to transform the complex original data into a sequence of discrete, quasi-stable meaningful patterns, also called states. It was shown that the distribution of these states may play a biomarker role in various psychiatric (Koenig et al., 1999; Janoos et al., 2013) and neurological conditions including MS (Gschwind et al., 2016). In the current work, however, a brain state is a typical configuration of spatial activity without actual knowledge of spatial coordinates of electrodes.
We earlier defined brain states, or cognetoms, as snapshots of repeatable patterns of highly correlated signals also understood as moments of distributed brain activity of a given time point (Janoos et al., 2011; Mórocz et al., 2012). In other words, states are defined as patterns of statistical dependencies or spatio-temporal correlations between brain regions measured across EEG electrodes. The states serve as lowest-level elements, or level-0 cognetoms, quasi as alphabet letters or building blocks to use in future aggregation cascade models of ever increasing cognetom complexity and abstraction level.
@@
In a second move, for the temporal coding of brain states, or brain states sequence, the best match distribution of brain states is detected for each time point along the EEG data time axis with weights shown for each brain state. Each time step in the original sequence is assigned to its nearest brain state cluster, producing a simplified temporal sequence of discrete brain state codes, each associated with signal intensities across frequency bands. In other words, our approach translates the multi-channel raw EEG signal of each subject into a sequence of brain states as qualitative feature with the associated weights as numerical quantitative feature.
@@ 3.2 Specific feature combinations for MS detection
The relative importance of 80 feature combinations (8 brain states x 5 frequency bands x 2 experimental conditions) that separate the two cohorts into healthy vs MS is outlined in figure 2. We observe that a set of feature combinations of high importance include both eyes-open and eyes-closed conditions and the delta frequency band that appear to drive the separation between healthy and MS subjects. Averaged frequency spectra for both cohorts are shown as graphs in C.
@@ 3.3 Disease predictor compared to clinical data
The brain states analysis technique separates with surprising ease and high probability healthy controls from patients with MS (figures 1 and 3). Considering the data to originate from single clinical-grade EEG recordings, we therefore examined additional objective neurological measures such as EDSS values or plain disease duration which we hypothesized will correlate with our EEG analysis. These two clinical measures arguably reflect disease related accumulation of neurological deficits.
@@
The plots in figure 3 imply a correlation of these two clinical measures with the brain states predictor space, which, we surmise, distinguishes the better the two cohorts, the more disease burden over time accumulates in the CNS. Patients with symptoms other than pure motor symptoms show a subtle trend with EDSS computed on 26 such cases (middle graph B). Regarding disease duration, however, all patients to the far left in the ‘control’ brain states predictor space (marked green with letters from a to g) feature a short disease duration of less than two years (mean 0.8 years) and cited motor and sensory deficits which may originate from lesions in the spinal cord or lower brain rather than upper brain parts. In contrast, patients to the right marked red, with a mean disease duration of 6.0 years, reported also vision, balance or brain stem disturbances beside motor and sensory issues in the limbs. Still, those patients in red to the right with a brief disease duration of 3 years or less, may in fact, we hypothesize, suffer from a longer, hidden disease duration, albeit clinically compensated, to amass enough supratentorial pathology that finally deranges their EEG patterns. We assume that an increase in lesion load in the cerebrum, the principal generator for the EEG signal, relates just vaguely and non-linearly to the emerging complex clinical neurological deficits, be that in cognition, EDSS, or then in form of clinical disease duration, or ‘true’ disease onset, an unknown entity for that matter.
@@
Figure 3: Distinction of control subjects from MS patients through brain states prediction. Disease predictor values correspond to the rotated coordinates as seen in Fig. 1B using principal components of the point distribution with the predictor value considered as primary principal component. (A) The upper graph illustrates this separability for both cohorts plotted against EDSS measures obtained in MS. Patients are shown as red or green dots and controls as blue. Seven outlier patients are found on the left marked with green characters (a-g) for further discussion under Results. (B) The middle graph plots separability against EDSS measures in MS patients alone. Only patients having symptoms other then motor symptoms were included. For these 26 patients a linear regression is shown as gray line. (C) The lower graph plots separability against disease duration in years. Six patients (including d,f) are excluded for no data on disease duration. Linear regression is shown as gray trend line. The source for the post-analysis correlation calculation is in the values for the brain state predictor space and EDSS or disease duration. The open-source python module linregress from scipy served for calculation (Virtanen et al., 2020).
@@ 4 Discussion
We successfully transformed a brain states hypothesis (Janoos et al., 2011) into a new analytic method for processing EEG data using a space-neutral quantitative embedding and modeling approach for brain states and applied the new technique to separate the two subject cohorts of an earlier EEG study about MS (Jouzizadeh et al., 2021). We observed several beneficial aspects when brain states analysis is added in a context as described here.
@@ 4.1 Implications
First we demonstrate with an initial implementation of our novel algorithm, a two-phase processing pipeline with brain states detection and brain states analysis, that already regular clinical EEG data clearly carry the necessary information to discriminate two cohorts as shown here. This is to our knowledge a first successful demonstration of such a hypothesis-free space-neutral data-driven full-brain raw-data EEG analysis approach for MS.
Second we state that the inclusion of brain states in the analysis is sensible and enhances the statistical robustness of the approach.
Third we believe this brain states approach is easily generalizable to other disease conditions and importantly also for the intra-individual monitoring of serially acquired longitudinal EEG data.
The current method readily distinguished healthy from MS cases using clinical EEG data in spite of obstacles faced with : i) early stage of the software implementation; ii) swap in the original purpose of the approach from monitoring serial data to cross-sectional comparing of subjects; iii) low information density in common clinical EEG recordings. We explain the analysis outcome with flow derangement between electrical mental processes in MS secondary to altered physical brain properties (Kisler et al., 2020; Paolicelli et al., 2021; Salim et al., 2021; Wątorek et al., 2024). This was newly confirmed in recent studies about disturbances of conductance velocity (Sorrentino et al., 2022; Wagner et al., 2025) although we currently offer no method to quantify processing speed in brain states. However, we already know that certain feature combinations of high relevance carry slow EEG components associated with both eye-open and eyes-closed conditions, that appear to prominently drive, but still as part of the entire train of feature combinations, the separation of controls from patients (see figure 2).
@@ 4.2 Significance
That this method succeeds on clinical-grade EEG data rests on several facts, in the processing of complete brain states, in the intrinsic noise suppression capacity, and in the characterizing of entire EEG recordings in form of comprehensive brain states sequences without cleansing noisy epochs. The need for artifact removal is an ongoing debate in the field while retaining the data in the most original state is a stringent reality (Delorme, 2023). In fact, the removal of artifactual, paradigm-related, non-brain-derived EEG components can detrimentally affect analysis of brain physiological data (McDermott et al., 2021), just as noise filtering may lead to undesired outcomes with poorly characterized noise patterns that are loosely related to behavior (Meisler et al., 2019). Experimental paradigms, however, come necessarily with random patterns of incidental and intentional muscular activities which surface as subject specific chance noise in space and time. The data processing pipeline during the brain state detection and the later analysis phases intrinsically suppress noise components (Janoos et al., 2011; Mórocz et al., 2012) and fundamentally render the approach apt to handle clinical-grade real-life data (Janoos et al., 2013). This ‘design-specific noise robustness’ lies in the fact that brain states are generated for the entire data set and hence hold inter-individual, universally present features but suppress intra-individual periodic or permanent signal properties. This way do patient-specific EEG artifacts due to motor, sensory and other effects not survive in brain states nor lead to subgroups in the analysis nor bias the subsequently computed disease predictor space.
Together these points highlight potential advantages of the method in that older less-advanced EEG databases could be sensibly re-analyzed with modern brain states analysis techniques, while collecting new EEG data in MS or other conditions could for budgetary reasons happen on affordable low-resolution EEG recording devices.
Having said that, only high sensor density (Formica et al., 2025; Wątorek et al., 2024), high scan frequency and rich cognitive stimulative task conditions will leverage dense enough information in the EEG and hence contribute to deep brain states abstraction models. While clinical EEG data come in their own right only the highest-quality EEG data can deliver the fodder to study the internal syntax inside multi-hierarchical cognetom cascade models.
@@ 4.3 Limitations
Third, the difference of mean age between the two cohorts amounts to nine years. We are, however, for the current retrospective analysis not aware of a correlation between age and disease severity (EDSS or symptoms) or disease duration. All subjects still belong to a young life stage where we expect task-less experiments to recruit similar brain functions and areas and so not to bias age-specifically the brain states predictor space.
Fourth, the brain states epoch window kept for data-technical reasons at two seconds width proved to work surprisingly well but does no justice to the speed at which ephemeral electrical brain activities operate. Hence this window size will be optimized in future iterations of this project.
Fifth, the current implementation of the BSD method is not suited for the topographic visualization of the 8 brain states, as it is in the case of the microstates method that follows a clear anatomical preponderance (Lehmann et al., 1987). Our current method focuses on a space-neutral approach for analyzing EEG data of patients with MS.
Sixth, we are unaware of the medication given to study participants nor was this information conveyed prior (Jouzizadeh et al., 2021). Certain drug classes are known to modify physiological brain responses where patients with MS may indeed receive periodically benzodiazepines (Van Schependom et al., 2019) for example to control muscle spasms and sleep disturbances, though such symptoms tend to occur at later disease stages. This makes a benzodiazepine induced bias in the analysis outcome of the present study less likely in view of the short mean disease duration of 5.1 years. Similarly does the predominant delta character of the leading brain states feature combinations (see figure 2) point away from the beta frequency presence in benzodiazepine treated patients with MS (Van Schependom et al., 2019).
@@ 4.4 Future objectives
First we see the purpose of this publication in the introduction of a novel hypothesis-free space-neutral data-driven analysis method tested on EEG and clinical data of patients with MS. However another reason is our desire to raise interest in the community towards functional brain data in longitudinal clinical studies in MS in the form of serial EEG measures (Kiiski et al., 2018).
We envision further improvements in the brain states detection (BSD) process through widening the spectrum of methods that analyze the EEG signal. For example the analysis through multifractality and persistence effectively characterizes EEG recordings of patients with MS during eyes open and eyes closed conditions (Wątorek et al., 2024). We consider multifractality and information about long-range temporal correlations, wavelet analysis and microstates, as further procedural candidates that enrich the brain states feature space and enhance model performance for time series analyses in future iterations of this project.